OpenStudy (anonymous):

I have to solve this in Algebra form: Absent minded Ralph leaves home at 8:00 AM for an important business meeting. Twenty minutes later, Mrs. Ralph discovers that Ralph has forgotten his briefcase. Ralph always drives 50 mph. If Mrs. Ralph drives 60 mph, what time will she catch up to Ralph?

OpenStudy (anonymous):

\[60mph(t) = 50mph(t) + (\frac{1}{3} \times 50mph)\]\[10t = \frac{50}{3}\]\[t = \frac{5}{3}h = 99.99 minutes\]

OpenStudy (anonymous):

he has gone one third of an hour so he has traveled \[\frac{50}{3}=16\tfrac{2}{3}\] miles already. then his distance is \[16\tfrac{2}{3}+50t\] and hers is \[60t\] and they meet at the same distance so solve \[\frac{50}{3}+50t=60t\] for t

OpenStudy (anonymous):

what alexray said